OpenStudy (anonymous):

The product of 4 and a number is 18 more than twice the number. What is the number?

OpenStudy (anonymous):

I almost had it...

OpenStudy (triciaal):

let the number = x product means multiply so 4* x 18 more than twice the number more than means add twice means 2times = 2*x put it all together 4 * x = 18 + 2 x now solve for x

OpenStudy (triciaal):

did you get the number is 9?

OpenStudy (rianiscool):

9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you @triciaal

OpenStudy (triciaal):

you are welcome.